Trying to help the children with homework. Cartesian Coordinate System. If I can see one problem finished I maybe able to understand.
Directions: Slope form both points
2nd point slope form and verify
Write in standard form
Label and intercept
1. (4,7) (8,3)

Answer by richwmiller(17219)

Cartesian coordinate system just means on graph paper where there is an x and y.
You aren't real clear what you want. But based on the words you used I assume you want the slope form and standard form of an equation that has those two points.
Something to remember : There are an infinite number of equations that have those points.
given :(4,7) and (8,3)
formulas
standard form ax+by+c=0
slope form y=mx+b the b's are not the same!
In y=mx+b m is the slope and b is the y intercept (y when x=0)
To find the slope, we use (y-y)/(x-x)= m the slope
The slope is the gradient or incline or slant of the line.
our y's are 7 and 3
our x's are 4 and 8
Just keep the first and second in their place.
7-3/4-8=4/-4=-1
our slope is -1 which means the line slants down from left to right.
Now we have y=-1x+b
y=-x+b
now pick either point and plug in the x and y
(4,7) x=4 y=7
7=-4+b
add 4 to both sides
11=b
now we have all the info needed.
y=-x+11 is the slope intercept form
when x=0 then y=11
Now to get the standard form in the form ax+by+c=0
We want everything on one side
y=-x+11
add x and subtract 11
y+x-11=0
almost done
we need x first so we rearrange.
x+y-11=0 is the standard form
I don't know what "Label and intercept" means.
Label what and intercept what?
Do you need a graph too? and need to label the intercepts?
I suspect you copied something incorrectly.